OpenStudy (anonymous):

The angle A is such that secA + tanA = 2 Show that secA-tanA=1/2 Hence, find the exact value of cosA.

OpenStudy (anonymous):

use (secA + tanA )(secA-tanA)=1 thus secA-tanA=1/2

OpenStudy (anonymous):

Solve simultaneously 2secA=5/2 cosA=4/5

OpenStudy (anonymous):

Don't really understand... Can you start from the beginning?

OpenStudy (anonymous):

You get the first part? right \[\sec^2A-\tan^2A=1=(secA-tanA)(secA+tanA)\]

OpenStudy (anonymous):

Not the "thus secA-tanA=1/2"

OpenStudy (anonymous):

Use result from first part with what is already given and add the two \[(secA+tanA)+(secA-tanA)=2+1/2\]

OpenStudy (anonymous):

\[\sec A + \tan A = \frac 1{(\sec A - \tan A)} \]

OpenStudy (anonymous):

Oh Ok.

OpenStudy (anonymous):

Note \(A \neq n\pi + \frac{\pi}{2} \) and \(n \in \mathbb{z} \)

OpenStudy (anonymous):

Ok, so how do you get the second part of the answer? @NotSObright and @FoolForMath

OpenStudy (anonymous):

Add the two equation.

OpenStudy (anonymous):

Already mentioned

OpenStudy (anonymous):

order order think think :)

OpenStudy (anonymous):

But how did you get this in the first place? 2secA=5/2 cosA=4/5 I'm trying!... I'm trying!.... It's difficult for me though.. :/

OpenStudy (anonymous):

sec A = 1/ cos A

OpenStudy (anonymous):

Still lost... :( My brain isn't working well today...

OpenStudy (anonymous):